OpenStudy (anonymous):

MathDude000 Juanita wants participate in 20 km bicycle race. She is considering using 1 of the 2 training programs below: Program A: Ride 10 km on the first day and increase the distance by an additional 500 m each day Program B: Ride 7,500 m on the first day and increase the distance by an additional 0.75 km each day. If Juanita chooses Program A, on which day she ride on the first 15 km for the first time

OpenStudy (whpalmer4):

With program A, the daily distance is 10,000 + 500(d-1) meters. With program B, the daily distance is 7,500 + 750(d-1) meters. The (d-1) is to account for the fact that we only add extra meters starting on the 2nd day. Take the expression representing daily distance for the plan and set it equal to 15,000. Solve for d.

OpenStudy (anonymous):

if juanita chooses program b how many total km will she ride in the first 3 days

OpenStudy (anonymous):

thats the equation i had trouble with

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

answer! cmon!!!!1

OpenStudy (whpalmer4):

Well, that wasn't part of what you asked, now is it ? :-) Okay, on day 1 with program b she rides 7500 + 750(d-1) = 7500 + 750(1-1) = 7500 On day 2 she rides 7500 + 750(2-1) = 7500 + 750(1) = 7500+750 = 8250 On day 3 she rides 7500 + 750(3-1) = (you do it) Add the 3 values together to get the distance ridden in the first 3 days

OpenStudy (anonymous):

So what would be part c

OpenStudy (whpalmer4):

Good question. I don't see anything labeled "part c" above, do you?

OpenStudy (anonymous):

There was

OpenStudy (anonymous):

Part c Juanita wants To ride at least 20 kilometers a day. Which program will allow her to reach this goal sooner.....??!???

OpenStudy (whpalmer4):

Do like I suggested in my initial post, except solve for 20,000 meters instead of 15,000. Compare the results you get, and choose the program which takes fewer days.

OpenStudy (anonymous):

then which program will allow her to reach this goal???????